OpenStudy (anonymous):

Prove: tan^2 θ•cos^2 θ + cos^2 θ = 1. You must show all work.

OpenStudy (anonymous):

\[\sin ^2/\cos ^2*\cos ^2+\cos ^2 = 1\]

OpenStudy (praxer):

tan^2 theta can be written as sin^2 theta / cos^2 theta so, sin^2 theta / cos^2 theta * cos^2 theta + cos^2theta sin^2 theta + cos^2 theta =1

OpenStudy (anonymous):

\[\sin ^2+\cos ^2 = 1 \]

OpenStudy (anonymous):

So you just cancel out the cos^2?

OpenStudy (anonymous):

\[1 = 1\] since we know that sin^2 +cos ^2 = 1

OpenStudy (anonymous):